OpenStudy (anonymous):

When do I use the trigonometric functions with the -1 on the top right of them that looks like an exponent? I've forgotten what they're called and I've forgotten when they're used.

OpenStudy (dape):

They are the inverse of the regular trigonometric functions, this means that for example \[ \sin(\sin^{-1}(x))=x \] and similarly for the other ones, they are used among other things for getting the angle when you know the ratio of the sides of a triangle. They are also called arcsin, arccos, arctan etc.

OpenStudy (anonymous):

So would I use the inverse if I didn't have the degree measure?

OpenStudy (dape):

Say you wanted to solve the equation \(\sin(\theta)=0.5\), then you could use arcsin on both sides to get \[ \arcsin(\sin(\theta))=\theta=\arcsin(0.5)=30^o \]

OpenStudy (dape):

Think of the inverses "cancelling" the other function.

OpenStudy (anonymous):

Oh, okay. I think I've got it now. Thank you for your help!